# Lifting Scheme-Based Implicit Disentanglement of Emotion-Related Facial Dynamics in the Wild

### v1

@@ Abstract
In-the-wild Dynamic facial expression recognition (DFER) encounters a significant challenge in recognizing emotion-related expressions, which are often temporally and spatially diluted by emotion-irrelevant expressions and global context respectively. Most of the prior DFER methods model tightly coupled spatiotemporal representations for direct classification, which may incorporate weakly relevant features such as facial contours and identity-specific characteristics, leading to information redundancy and emotion-irrelevant context bias. Several DFER methods have highlighted the significance of dynamic information for DFER, but utilize explicit manners to extract dynamic features with overly strong prior knowledge. In this paper, we propose a novel Implicit Facial Dynamics Disentanglement framework (IFDD). Through expanding wavelet lifting scheme to fully learnable framework, IFDD disentangles emotion-related dynamic information from emotion-irrelevant global context in an implicit manner, i.e., without exploit operations and external guidance. The disentanglement process of IFDD contains two stages, i.e., Inter-frame Static-dynamic Splitting Module (ISSM) for rough disentanglement estimation and Lifting-based Aggregation-Disentanglement Module (LADM) for further refinement. Specifically, ISSM explores inter-frame correlation to generate content-aware splitting indices on-the-fly. We preliminarily utilize these indices to split frame features into two groups, one with greater global similarity, and the other with more unique dynamic features. Subsequently, LADM first aggregates these two groups of features to obtain fine-grained global context features by an updater, and then disentangles emotion-related facial dynamic features from the global context by a predictor. Extensive experiments on in-the-wild datasets have demonstrated that IFDD outperforms prior supervised DFER methods with higher recognition accuracy and comparable efficiency. Code is available at https://github.com/CyberPegasus/IFDD.

### v2

@@ Abstract
In-the-wild dynamic facial expression recognition (DFER) encounters a significant challenge in recognizing emotion-related expressions, which are often temporally and spatially diluted by emotion-irrelevant expressions and global context. Most prior DFER methods directly utilize coupled spatiotemporal representations that may incorporate weakly relevant features with emotion-irrelevant context bias. Several DFER methods highlight dynamic information for DFER, but following explicit guidance that may be vulnerable to irrelevant motion. In this paper, we propose a novel Implicit Facial Dynamics Disentanglement framework (IFDD). Through expanding wavelet lifting scheme to fully learnable framework, IFDD disentangles emotion-related dynamic information from emotion-irrelevant global context in an implicit manner, i.e., without exploit operations and external guidance. The disentanglement process contains two stages. The first is Inter-frame Static-dynamic Splitting Module (ISSM) for rough disentanglement estimation, which explores inter-frame correlation to generate content-aware splitting indexes on-the-fly. We utilize these indexes to split frame features into two groups, one with greater global similarity, and the other with more unique dynamic features. The second stage is Lifting-based Aggregation-Disentanglement Module (LADM) for further refinement. LADM first aggregates two groups of features from ISSM to obtain fine-grained global context features by an updater, and then disentangles emotion-related facial dynamic features from the global context by a predictor. Extensive experiments on in-the-wild datasets have demonstrated that IFDD outperforms prior supervised DFER methods with higher recognition accuracy and comparable efficiency. Code is available at https://github.com/CyberPegasus/IFDD.
@@ Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (grant U2441244), and by Zhejiang Provincial Natural Science Foundation of China (grant LZ24F030006).